# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v1

@@ 4.4 Ablation Studies
Table 6: Ablation for obtaining the final performance.
Table 6 provides ablation experiments to obtain the final performance. First, by increasing the training time from 15 epochs to 24 epochs, an improvement of 0.57 mIoU can be achieved. Then, by using CBGS [61], another improvement of 0.43 mIoU can be achieved. When using CBGS, since resampling increases the number of samples in each epoch, for the sake of a fair comparison, we limited the training to 6 epochs to approximate the total training time of 24 epochs without CBGS.
4.5 Visualization
Figure 3 presents the visualizations of DAOcc on Occ3D-nuScenes validation set without using camera mask. The results show that our DAOcc made a relatively complete prediction of the scene, accurately reproducing it in fine detail.
5 Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, aiming to achieve superior performance while using a deployment-friendly image feature extraction network and input image resolution. Extensive experiments on Occ3D-nuScenes and SurroundOcc demonstrate the superiority of our approach over existing methods. We believe that DAOcc can be put into practice and inspire future research.

### v2

@@ 8 Detection Performance
Table 8: Ablation for achieving final performance. EP15, EP24, and EP6 represent training for 15 epochs, 24 epochs, and 6 epochs, respectively.
9 Visualization
Figure 4 presents the visualizations of DAOcc on Occ3D-nuScenes validation set. The camera visible mask is not used during model training and visualization. The results demonstrate that DAOcc is capable of making a relatively complete prediction of the scene, accurately reproducing it in fine detail.
Dataset Camera Mask mIoU mAP NDS Car Truck Bus T.L. C.V. Ped. M.C. B.C. T.C. B.R.
Occ3D-nuScenes [44] 53.8 59.4 64.3 85.7 57.2 65.8 33.1 22.8 75.5 68.3 47.3 70.2 68.2
Occ3D-nuScenes [44] 47.9 59.6 64.2 85.4 55.0 64.6 34.1 23.1 76.3 67.0 50.7 70.8 69.0
SurroundOcc [51] - 30.5 53.6 60.2 81.9 42.7 56.0 26.3 19.7 72.6 57.7 43.8 66.7 68.1
Table 9: 3D occupancy prediction and detection performance of DAOcc. ‘T.L.’, ‘C.V.’, ‘Ped.’, ‘M.C.’, ‘B.C.’, ‘T.C.’, and ‘B.R.’ are short for trailer, construction vehicle, pedestrian, motorcycle, bicycle, traffic cone and barrier, respectively.
Figure 4: Qualitative visualizations on Occ3D-nuScenes validation set. The camera visible mask is not used during model training and visualization.